### 07 Apr 2026 v2

# Spectral study of the pseudogap in unitary Fermi gases

Abstract
The recent observation of a pseudogap in unitary Fermi gases [Li et al., Nature 626, 288 (2024)] provides strong evidence for a pairing origin of the pseudogap in Fermi superfluids. Here we present a quantitative spectral study of these gases, comparing our theoretically calculated momentum‑resolved rf/microwave spectra with the experimental data. Using an iterative treatment of the fermion self‑energy beyond the previous pseudogap approximation, based on a pairing fluctuation theory that incorporates both particle‑particle and particle‑hole $T$ matrices with self‑consistent feedback, we achieve excellent agreement with the measured spectral functions and extracted pseudogap parameters. Our results not only explain the experimental data microscopically but also reinforce the pairing origin for the pseudogap and the pairing‑fluctuation theory of Fermi superfluidity.
Atomic Fermi gases have provided an ideal platform for simulating and studying the physics of Fermi superfluidity and superconductivity in solids [4], such as high $T_{c}$ superconductivity in cuprates. One central issue that perplexes our understanding of the high $T_{c}$ superconductivity is the very existence of a pseudogap in the single-particle excitation spectrum in the optimally doped and underdoped cuprates [25]. Assuming a pairing origin for the pseudogap, as advocated in the preformed pair scenario among competing theories for high $T_{c}$ superconductivity, one would expect to see a pseudogap in atomic Fermi gases when the pairing interaction is tuned to be strong. Indeed, this was proposed about the same time [23] when Fermi superfluidity was realized experimentally in these gases two decades ago [20]
The pairing and superfluid phenomena in ultracold atomic Fermi gases, particularly in the unitary regime, have been extensively studied over the past two decades. A key focus is the pseudogap phenomenon [4, 5, 16]; the existence of a pseudogap in strongly interacting Fermi gases will demonstrate that a strong pairing interaction can lead to the formation of a pseudogap [4], and thus provide naturally a plausible explanation for the widespread pseudogap phenomena in high $T_{c}$ superconductors [25]. Conclusive evidence came from very recent high-precision momentum-resolved microwave spectroscopy measurements in a homogeneous unitary Fermi gas of 6Li [14]. It is the purpose of this work to provide microscopic calculations that are in quantitative agreement with the experiment, thereby strengthening the support for both the pairing origin of the pseudogap and our particular pairing fluctuation theory [2].
Pseudogap can be directly probed via measuring the fermion spectral function using, e.g., angle-resolved photoemission spectroscopy (ARPES) in solid state superconductors [9] or its counterpart, momentum-resolved rf or microwave spectroscopy, in atomic Fermi gases [24, 1]. Indeed, the existence of the pseudogap in Fermi gases has been experimentally investigated in a trap using either momentum integrated [8] or momentum-resolved rf spectroscopy [24]. However, these earlier rf measurements were plagued by trap inhomogeneity, limited resolution, low signal-to-noise ratio, and final-state interactions [19, 10], and thus have allowed room for ambiguity in explaining the rf spectra, with possible alternative interpretations without invoking fermion pairing [21, 17]. The measurements of Li et al [14] were free of the complications caused by the trap inhomogeneity and the well-known final-state effects [22]. The evidence for pseudogap includes clear, simultaneous observation of the two BCS-like branches of the quasiparticle dispersion in a homogeneous unitary Fermi gas of 6Li both below and above the superfluid transition temperature $T_{c}$, and quantitative measurements of the excitation gap and Cooper pair decay rate as a function of temperature.
In this Letter, we consider a homogeneous three-dimensional (3D) ultracold Fermi gas at unitarity with short-range $s$-wave pairing interaction $V_{\mathbf{k},\mathbf{k}^{\prime}}=g<0$, which provides a good description of the 6Li gases studied in Ref. [14]. We use a pairing fluctuation theory [7] that incorporates particle-hole (ph) channel effects in the particle-particle scattering $T$ matrix. To get quantitatively accurate results of the spectral function, we go beyond previous approximations for the pseudogap self-energy and treat it with elaborate iteration with full numerics. We arrive at a spectral function $A(\mathbf{k},\omega)$ that is in quantitative agreement with experimental data, with two BCS-like quasiparticle branches in the 3D spectral intensity map and the right amount of broadening. Treating these theoretically generated data with the same experimental procedure, we fit the loci of the spectral intensity peaks with BCS-like dispersions, and the energy distribution curves (EDCs) with the same phenomenological model for the self-energy. The extracted pairing (pseudo)gap, Hartree energy $U$, and effective fermion mass $m^{*}$, as well as the inverse pair lifetime $\Gamma_{0}$ and single-particle scattering rate $\Gamma_{1}$, are in quantitative agreement with experiment, both below and above $T_{\text{c}}$. This also includes the thermally activated exponential behavior of $\Gamma_{0}$ and nearly constant behavior of $\Gamma_{1}$ versus $T$. Our results not only provide a quantitative explanation for the experiment, but also manifest that our theory captures the right physics.
To be concrete, the fermion self-energy in our theory [4, 2] contains two terms, associated with the contributions of the Cooper pair condensate and finite-momentum pairing fluctuations, given by $\Sigma_{\text{sc}}(K)=-\Delta_{\text{sc}}^{2}G_{0}(-K)$ and $\Sigma_{\text{pg}}(K)=\sum_{Q\neq 0}t_{\text{pg}}(Q)G_{0}(Q-K)$, respectively. Here the superfluid order parameter $\Delta_{\text{sc}}$ vanishes at $T\geq T_{c}$, and $G_{0}(K)=1/(\mathrm{i}\omega-\xi_{\mathbf{k}})$ is the bare Green’s function, with the free fermion dispersion $\xi_{\mathbf{k}}\equiv\epsilon_{\mathbf{k}}-\mu^{\prime}=k^{2}/2m-\mu^{\prime}$ and chemical potential $\mu^{\prime}$, which contains a shift due to the Hartree energy (See Equation (2.4) of Ref. [6]). Here the Hartree shift is generalized to include the whole diagonal part of the self-energy at the Fermi level, so that $G_{0}^{-1}(0,k_{\mu})=G^{-1}(0,k_{\mu})=0$. The $T$-matrix associated with finite-momentum pairs, $t_{\text{pg}}(Q)=1/(g^{-1}+\chi(Q))$, is given by an infinite series of particle-particle scattering ladder diagrams, with pair susceptibility $\chi(Q)=\sum_{K}G_{0}(Q-K)G(K)$, where $G(K)$ is the full Green’s function. Note that this mixed form of $\chi(Q)$, derived from the equations of motion approach [11, 6], distinguishes this pairing fluctuation theory from other $T$-matrix-based theories [12]. Now the total self-energy is given by
$\Sigma(K)$ $=$ $\Sigma_{\text{sc}}(K)+\Sigma_{\text{pg}}(K)$ (1)
$=$ $-\Delta^{2}_{\text{sc}}G_{0}(-K)+\sum_{Q\neq 0}t_{\text{pg}}(Q)G_{0}(Q-K)\,.$
Here and throughout, we set the volume to unity, take the natural units $\hbar=k_{\text{B}}=1$, and use the four-momentum notation $K\equiv(\mathbf{k},\mathrm{i}\omega_{n})$, $Q\equiv(\mathbf{q},\mathrm{i}\Omega_{l})$, $\sum_{Q}\equiv T\sum_{l}\sum_{\mathbf{q}}$, etc, where $\omega_{n}$ (fermionic) and $\Omega_{l}$ (bosonic) are Matsubara frequencies [2].
The Thouless criterion for superfluidity gives $t^{-1}_{\text{pg}}(0)=0$. This leads to the pseudogap approximation $\Sigma_{\text{pg}}(K)\approx-\Delta_{\text{pg}}^{2}G_{0}(-K)$ for $T\leq T_{c}$ and also above $T_{c}$, as long as the pair chemical potential $\mu_{\text{p}}$ is small. This defines the pseudogap parameter via $\Delta_{\text{pg}}^{2}=-\sum_{Q}t_{\text{pg}}(Q)$. It brings both $\Sigma(K)$ and $G(K)$ into the simple BCS-like form, with $\Sigma(K)=-\Delta^{2}G_{0}(-K)$ and $G(K)={u_{\mathbf{k}}^{2}}/(\mathrm{i}\omega_{n}-E_{\mathbf{k}})+{v_{\mathbf{k}}^{2}}/(\mathrm{i}\omega_{n}+E_{\mathbf{k}})$, where the coherence factors are $u_{\mathbf{k}}^{2}=(1+\xi_{\mathbf{k}}/E_{\mathbf{k}})/2$, $v_{\mathbf{k}}^{2}=(1-\xi_{\mathbf{k}}/E_{\mathbf{k}})/2$, with $E_{\mathbf{k}}=\sqrt{\xi_{\mathbf{k}}^{2}+\Delta^{2}}$, and the pairing gap $\Delta=\sqrt{\Delta^{2}_{\text{sc}}+\Delta^{2}_{\text{pg}}}$. This BCS-like $G(K)$ then yields the spectral function
$A_{\text{init}}(\mathbf{k},\omega)=2\pi\left[u_{\mathbf{k}}^{2}\delta(\omega-E_{\mathbf{k}})+v_{\mathbf{k}}^{2}\delta(\omega+E_{\mathbf{k}})\right]\,$, (2)
which will be used as the initial input in our iterative numerical procedure described below.
While the pseudogap approximation makes it easy to solve for $T_{c}$ and the gap parameters, it is an oversimplification. It drops the diagonal part of the self-energy so that $\Sigma_{\text{pg}}$ vanishes when $\Delta_{\text{pg}}=0$. The resulting spectral function is obviously too sharp compared with that one would expect from the convoluted form of $\Sigma_{\text{pg}}$; The finite-momentum pair distribution will spread out the spectral weight.
To enable quantitative comparisons with experiment [14], we need to include the contribution of particle-hole fluctuations as well. Following Ref. [7], we incorporate the entire particle-hole $T$ matrix $t_{\text{ph}}$ such that it serves as a screened pairing interaction, and causes a non-uniform shift of the inverse pairing interaction, $1/g$, by the particle-hole susceptibility $\langle\chi_{\text{ph}}\rangle$ averaged near the Fermi surface. Hence the finite-momentum $T$-matrix becomes $t_{\text{pg}}(Q)=1/(g^{-1}+\chi(Q)+\langle\chi_{\text{ph}}\rangle)$. With the BCS-like $G(K)$, $t^{-1}_{\text{pg}}(Q)$ is given by
$t^{-1}_{\text{pg}}({\mathbf{q}},\mathrm{i}\Omega_{n})\!$ $=$ $\!\sum_{\mathbf{k}}\left[\frac{1-f(E_{\mathbf{k}})-f(\epsilon_{\mathbf{k-q}})}{E_{\mathbf{k}}+\epsilon_{\mathbf{k-q}}-\mathrm{i}\Omega_{n}}u_{\mathbf{k}}^{2}\right$. (3)
- $\!\left.\frac{f(E_{\mathbf{k}})-f(\epsilon_{\mathbf{k-q}})}{E_{\mathbf{k}}-\epsilon_{\mathbf{k-q}}+\mathrm{i}\Omega_{n}}v_{\mathbf{k}}^{2}\right]+\langle\chi_{\text{ph}}\rangle+\frac{1}{g}\,.$
Furthermore, rather than using the pseudogap approximation, we calculate $\Sigma_{\text{pg}}(K)$ directly using the full convolution in Eq. (1), which leads to
$\Sigma_{\text{pg}}^{\text{R}}(\mathbf{k},\omega)=-\!\sum_{\mathbf{q}}\!\int\!{\frac{{\mathrm{d}}\Omega}{\pi}}{\frac{b(\Omega)+f(\xi_{\mathbf{q}-\mathbf{k}})}{\xi_{\mathbf{q}-\mathbf{k}}\!-\Omega+\omega+\mathrm{i}0^{+}}}\,\text{Im}\,t_{\text{pg}}^{\text{R}}(\mathbf{q},\Omega)\,$, (4)
after analytical continuation. Here $b(x)$ and $f(x)$ are Bose and Fermi distribution functions, respectively. The $b(x)$ term describes a dominantly off-diagonal BCS-like self-energy due to pair formation. The $f(x)$ term, which is dropped in the pseudogap approximation, reflects the interaction effect of a fermion with the particle-hole bubbles from the medium, and has a regular, dominantly “diagonal” contribution to the Hartree self-energy. The imaginary part of the self-energy thus reflects the finite lifetimes of quasiparticles associated with each process.
Figure 1: (a) Real and (b) imaginary parts of the retarded self-energy $\Sigma^{\text{R}}(k_{\mu},\omega)$ for a unitary Fermi gas at $|\mathbf{k}|=k_{\mu}$ and $T_{\text{c}}$. Red dashed and green dash-dotted lines show the Bose and Fermi components of total self-energy (blue solid lines), respectively, along with the average Hartree energy $\bar{E}_{\text{Hartree}}$ (cyan dotted line).
Using the Kramers-Kronig relation, we obtain the real and imaginary parts of $\Sigma_{\text{pg}}^{\text{R}}(\mathbf{k},\omega)$, as shown in Fig. 1. Note that the off diagonal self-energy must necessarily vanish at $k=k_{\mu}$ and $\omega=0$, where $k_{\mu}=\sqrt{2m\mu^{\prime}}$ corresponds to the wave vector on the Fermi surface, where the back-bending of the BCS-like quasiparticle dispersion occurs. This requires that $\bar{E}_{\text{Hartree}}={\text{Re}}\Sigma^{\text{R}}_{\text{pg}}(k_{\mu},0)$. Thus the physical chemical potential is given by $\mu=\mu^{\prime}+\bar{E}_{\text{Hartree}}$. Note that the main effect of the Hartree-dominated diagonal self-energy is a chemical potential shift and slight fermion mass renormalization (It should now be clear that our bare Green’s function $G_{0}(K)$ contains the diagonal self-energy via the shifted chemical potential $\mu^{\prime}$.) Now the spectral function is
$A(\mathbf{k},\omega)=\frac{-2\,{\text{Im}}\Sigma^{\text{R}}(\mathbf{k},\omega)}{[\omega-\epsilon_{\mathbf{k}}+\mu-\text{Re}\Sigma^{\text{R}}(\mathbf{k},\omega)]^{2}+[{\text{Im}}\Sigma^{\text{R}}(\mathbf{k},\omega)]^{2}}\,.$ (5)
The pseudogap in the spectral function originates from the negative peak around $\omega=0$ in $\text{Im}\Sigma^{\text{R}}(k_{\mu},\omega)$, where $\Sigma^{\text{R}}=\Sigma_{\text{sc}}^{\text{R}}+\Sigma_{\text{pg}}^{\text{R}}$. Detailed derivations of our iterative framework and quantitative numerical results in the BCS and BEC regimes are given in a companion paper [13].
Figure 2: Contour plot of $\mathbf{k}^{2}A(\mathbf{k},\omega)$ at (a) $T/T_{\text{c}}=0.77$, (b) 11, (c) $1.11$, and (d) $1.51$, with $T_{\text{c}}/T_{\text{F}}=0.2$, showing quasiparticle dispersions evolving from BCS-like gapped branches to a single S-shaped branch with a decreasing $\Delta$.
Shown in Fig. 2 are contour plots of the spectral intensity $\mathbf{k}^{2}A(\mathbf{k},\omega)$ as a function of $k=|\mathbf{k}|$ and $\omega$ at (a) $T/T_{\text{c}}=0.77$, (b) 11, (c) $1.11$ and (d) $1.51$, matching temperatures in Ref. [14], with $T_{\text{c}}/T_{\text{F}}=0.2$. At $T/T_{\text{c}}\leq 1$ in panels (a) and (b), two clearly resolved excitation branches, with a sizable pairing gap $\Delta$, correspond to the particle- and hole-like dispersions. Both branches exhibit a back-bending behavior around $k_{\mu}\simeq 0.93k_{\text{F}}$, characteristic of BCS-like dispersions. As temperature rises from panel (b) to (c), $\Delta$ shrinks, resulting in a hybridization between the two branches and the emergence of an $S$-shaped dispersion, as observed in Ref. [14]. Intuitively, at higher $T$ with smaller gap, the spectral weight of the low $k$ part of the upper branch and the large $k$ part of the lower branch is expected to become smaller with a wider spread in frequency. This renders the spectral peak essentially invisible in these two parts, hence leaving an $S$-shaped dispersion. This also indicates less well-defined quasiparticles at this temperature above $T_{\text{c}}$, so that the dispersion, as given by the locus of the spectral peak, evolves continuously from the lower to the upper branch. At even higher $T/T_{\text{c}}=1.51$ in panel (d), $\Delta$ becomes so small that the $S$ shape is invisible and the dispersion looks simply parabolic.
Figure 3: Overlay of fitted dispersions on top of the spectral intensity map of $\mathbf{k}^{2}A(\mathbf{k},\omega)$ for (a) $T/T_{\text{c}}=0.9$ and (b) $1.11$. Red dots indicate $E_{\text{max}}(k)$; orange and blue lines correspond to $E_{k}^{\pm}$, respectively. Comparison between theory (blue squares) and experiment (orange triangles) for (c) $\Delta$ and (d) $U$, as well as (inset) $m^{*}$, at different $T/T_{\text{c}}$. Error bars represent one standard deviation.
To compare with experiment, we need to extract the gap from the calculated spectral data, rather than using the initial input in Eq. (2), following the same experimental procedure of data analysis as in Ref. [14]. We fit with BCS-like dispersions the locus of the maximum spectral response, $E_{\text{max}}(k)$, in $\mathbf{k}^{2}A(\mathbf{k},\omega)$. The particle- and hole-like dispersions are given by
$E_{{k}}^{(\pm)}=\pm\sqrt{\left(\frac{k^{2}}{2m^{*}}+U-\mu\right)^{2}+\Delta^{2}}\,$, (6)
where $U$ is the chemical potential shift, and $m^{*}$ is the renormalized fermion mass, both arising from the Hartree energy. The $S$-shaped dispersion is fitted with a weighted combination,
$E_{k}=a_{k}^{2}E_{k}^{(+)}+b_{k}^{2}E_{k}^{(-)}\,$, (7)
where $a_{k}^{2},b_{k}^{2}=\frac{1}{2}[1\pm\tanh(\frac{k-k_{\text{c}}}{\sigma})]$, with $k_{\text{c}}$ and $\sigma$ representing the crossover momentum and the crossover rate between the two branches, respectively.
In Fig. 3, we present the fitted dispersion $E_{{k}}^{(\pm)}$ (orange and blue lines) overlaid on top of the spectral intensity map for (a) $T/T_{\text{c}}=0.9$ and (b) $1.11$ using Eq. (6), along with an $S$-shape fit (green line) in panel (b) using Eq. (7). Panels (c) and (d) show the fitting parameters $\Delta$, $U$, and $m^{*}$ (inset) for the hole-like branch as a function of $T/T_{\text{c}}$. We use $E_{k}^{(-)}$ in Eq. (6) to fit the lower branch for $0.77\leq T/T_{\text{c}}\leq 1.04$, and $E_{k}$ in Eq. (7) to fit S-shaped dispersions for $1.11\leq T/T_{\text{c}}<1.51$. For $T/T_{\text{c}}=1.51$, a parabolic fit with $E_{k}={k^{2}}/{2m^{*}}+U-\mu$ yields a lower error than the S-shape fit. The comparison of our theoretical values (blue squares) with the corresponding experimental results (orange triangles) from Ref. [14] shows a quantitative agreement, despite a slight deviation in $U$ above $T_{\text{c}}$. This deviation may have to do with the fact that the numerical iteration is done only once. It should also be noted that fitting the upper branch yields slightly different results. Such particle-hole asymmetry becomes more pronounced with increasing interaction strength beyond unitarity [13].
Figure 4: EDCs at (a) $T/T_{\text{c}}=1$ and (b) $T/T_{\text{c}}=1.11$ for various $k/k_{\text{F}}$ near $k_{\mu}$, showing the quasiparticle peak evolution.
Next, shown in Fig. 4 is the evolution of the EDCs of $A(\mathbf{k},\omega)$ for a series of $k/k_{\text{F}}$ not far from $k_{\mu}$ at (a) $T/T_{\text{c}}=1$ and (b) $1.11$. At $T_{c}$ (Fig. 4(a)), the EDC is composed of two clearly resolved spectral peaks, reflecting well-defined particle- and hole-like quasiparticle energies. As $k/k_{\text{F}}$ increases, the spectral weight gradually shifts from the hole-like to the particle-like branch, as described by the coherence factors $v_{\mathbf{k}}^{2}$ and $u_{\mathbf{k}}^{2}$. In contrast, for the above $T_{c}$ case in panel (b), the EDCs exhibit only a single peak; here the BCS-like quasiparticles are no longer sharply defined. Nevertheless, the strong asymmetry of the peaks and the large peak width at $k/k_{F}=0.95$ do suggest that this single peak results from merging two broad peaks rather than closing the pairing gap.
Figure 5: (a) Normalized EDCs of $A(\mathbf{k},\omega)$ from our calculations (blue) at $k=0.93k_{\text{F}}$ for different $T/T_{\text{c}}$, as labeled, (b) numerical $\Delta$ (blue) (c) $\Gamma_{0}$ and $\Gamma_{1}$ (inset) from the EDC fit. For comparison, the corresponding experimental data are also shown (in orange). The theoretical $\Gamma_{0}$ data are also fitted with an exponentially activated behavior (blue line) in (c). Error bars represent one standard deviation.
Finally, we analyze the EDCs at the back-bending point $k_{\mu}=0.93k_{\text{F}}$ (whose temperature dependence we ignore, following the experiment [14]). We also apply a Gaussian broadening with a fixed standard deviation of $0.16E_{\text{F}}$ when calculating the EDCs, to account for instrumental resolution [14] which reflects the overall error caused by the uncertainty in both frequency and momentum measurements. In Fig. 5(a), we show our theoretical normalized EDCs (blue squares) and compare with experimental data (orange circles) at $k=0.93k_{\text{F}}$ for $T/T_{\text{c}}$ ranging from $0.77$ to $1.51$. With increasing $T$, the quasiparticle peaks gradually broaden and merge, indicating a decreasing $\Delta$, and thus a disappearing peak-to-peak separation. The double peak structure becomes a single broad peak as $T$ surpasses $T_{\text{c}}$, even though the fitted $\Delta$ is still nonzero. The calculated EDCs match experimental data closely, with a slight discrepancy in the peak height of the hole-like branch near $T_{\text{c}}$. This primarily arises from a slight deviation between the actual $k_{\mu}$ and the assumed $k_{\mu}=0.93k_{\text{F}}$, which highlights the sensitivity of EDCs to $k/k_{F}$ relative to the Fermi level, as can be seen in Fig. 4.
We fit the EDCs using the same phenomenological self-energy model [15, 18, 3] as in Ref. [14], given by
$\Sigma(\mathbf{k},\omega)=\frac{\Delta^{2}}{\omega+\xi^{*}(\mathbf{k})+\mathrm{i}\Gamma_{0}}-{\mathrm{i}}\Gamma_{1}+[\xi^{*}(\mathbf{k})-\epsilon_{\mathbf{k}}+\mu]\,$, (8)
where $\xi^{*}(\mathbf{k})\equiv\mathbf{k}^{2}/2m^{*}+U-\mu$ satisfying $\xi^{*}(k_{\mu})=0$, $\Gamma_{0}$ denotes the inverse pair lifetime, and $\Gamma_{1}$ represents the $\omega$-independent single-particle scattering rate. The expression of the spectral function derived from this model via Eq. (5) is used to fit the EDCs to extract $\Delta$, $\Gamma_{0}$, and $\Gamma_{1}$. The results are shown in Fig. 5(b) and Fig. 5(c), along with the experimental data for quantitative comparison. The agreement is good. In particular, we have a finite pseudogap at and above $T_{c}$, with $\Delta\approx 0.3E_{\text{F}}$ at $T_{c}$. As shown in Fig. 5(c), the theoretical data points of $\Gamma_{0}$ (blue squares) fit nicely with an exponential, thermally activated behavior, $\Gamma_{0}\propto\exp(-2\Delta_{\text{0}}/T)$ (blue curve), suggesting that the inverse pair lifetime is governed by the virtual pair breaking and recombination process, with excitation energy $2\Delta_{0}$. The fit yields $\Delta_{0}/E_{\text{F}}=0.41$, slightly larger than the pairing gap at $T=0.65T_{\text{c}}$, the lowest temperature accessed in experiment. The single-particle scattering rate $\Gamma_{1}$ shown in the inset of Fig. 5(c) remains nearly temperature independent, insensitive to pairing. Performing Gaussian broadening on the EDCs mainly leads to an overall increase in $\Gamma_{0}$ (and $\Gamma_{1}$) even at zero $T$, which is removed by background subtraction.
In summary, we have performed a spectral study of the pseudogap phenomena in unitary Fermi gases. By going beyond the previous pseudogap approximation for the self-energy, and incorporating the particle-hole channel screening to the interaction, we have arrived at a quantitative agreement with the recent experiment [14] on the spectral function, and the associated pairing gap $\Delta$, inverse pair lifetime $\Gamma_{0}$, and single-particle scattering rate $\Gamma_{1}$, using the same data analysis procedure as in Ref. [14]. In particular, the pseudogap self-energy is calculated numerically using full convolution, which captures explicitly the Hartree shift previously dropped (or hidden in chemical potential) in the pseudogap approximation. Our work thus provides a quantitative microscopic explanation for the observed pseudogap and establishes pairing fluctuations as the key mechanism underlying the pseudogap in unitary Fermi gases.
We thank Xi Li, Xing-Can Yao and Zhiqiang Wang for useful discussions. This work was supported by the Quantum Science and Technology - National Science and Technology Major Project (Grant No. 2021ZD0301904).